# DeepFleet: Multi-Agent Foundation Models for Mobile Robots

### v1

@@ 1 Introduction
The discovery that scaling up deep neural networks and their training data can unlock unprecedented capabilities in natural language understanding  [1], speech recognition  [2], vision [3], and reasoning [4] has recently led to efforts to apply the same data-driven scaling approach to robotic systems [5, 6, 7, 8]. Inspired by these efforts, we explore the design space of multi-agent foundation models to support the development of the next generation of robot fleet coordination algorithms. Just as foundation models pretrained on simple tasks such as masked text prediction can be adapted to applications via supervised fine-tuning, reinforcement learning, knowledge distillation, or in-context learning, our multi-agent fleet models are designed to serve as a foundation for many applications in our warehouses, such as congestion forecasting, adaptive routing, and proactive rescheduling.
@@ 2 Related Work
Significant work has been dedicated to modeling multi-agent interactions in trajectory forecasting, particularly in domains such as pedestrian motion prediction, autonomous driving, and social navigation. Early works such as Social-LSTM [9] and Social-GAN [10] captured inter-agent dynamics via recurrent neural networks with interaction-aware pooling. Transformer-based architectures [11, 12], variational models [13, 14], and hybrid methods combining GNNs with probabilistic forecasting [15, 16] have since improved generalization and multimodal prediction capabilities. However, these models are primarily applied to small-scale environments with limited agent counts and are not designed to scale to the thousands of agents encountered in warehouse scenarios.
@@
Research in warehouse robotics has historically focused on task assignment, multi-agent pathfinding (MAPF), and traffic management. Classic heuristic-based methods such as conflict-based search (CBS) [17] and prioritized planning [18] provide strong guarantees but struggle to scale in dense, real-time environments. Learning-based alternatives have emerged to address path planning and task reallocation [19, 20], but they typically lack generalization to diverse layouts and fail to incorporate long-horizon prediction. These methods are often designed for specific optimization objectives rather than as general-purpose predictive models. Recent work [21] introduces scalable imitation learning for lifelong multi-agent path finding, enabling robust coordination of up to 10,000 robots by combining efficient communication and global guidance to outperform both learning- and search-based baselines. Similarly, MAPF-GPT [22] leverages transformer-based imitation learning to efficiently solve large-scale pathfinding problems, demonstrating strong generalization across diverse environments. While these learning-based approaches rely on simulated data, we use a unique large-scale dataset of real warehouse operations.
@@
Recent work in spatiotemporal modeling has produced powerful neural representations for structured environments. TrajectoryCNN [23], Scene-Transformer [24], and ST-GNNs [25] illustrate the utility of combining convolutional or graph-based spatial reasoning with temporal sequence modeling. In robotics, structured state-space models (SSMs) [26, 27] and graph-based encodings [28] have improved data efficiency and temporal coherence. Notably, GraphCast [29] and GenCast [30] have demonstrated the effectiveness of graph-based and generative models for large-scale spatiotemporal prediction, respectively excelling at deterministic and probabilistic forecasting in highly complex systems. Nonetheless, these methods often lack scalability, task generalization, or the integration of diverse spatial and temporal contexts required for warehouse-scale fleet modeling.
@@
The rise of foundation models was enabled by empirical studies [31] that demonstrated that language model performance followed predictable scaling laws with respect to compute, model size, and data. After further refinement [32], these principles have been used to inform the scaling of large multi-modal models such as PaLM [33] and LLaMA [34].
@@
Foundation models are also achieving state-of-the-art capabilities in robotics. Models such as RT-1 [35], RT-2 [36], VIMA [37], and PerAct [38] have demonstrated strong performance in manipulation and visuomotor policy learning. Flamingo [39] has enabled visual reasoning for embodied agents, while $\pi_{0}$ [6], Octo [8], and Helix [40] have extended this foundation model approach toward multi-task, multi-embodiment settings. DeepMind’s work on robotics foundation models and scalable policy architectures [41] highlights the increasing effort to unify policy learning, perception, and reasoning in a single framework. These models emphasize generalization, multi-task capability, and real-world robustness—but are still primarily focused on manipulation or low-agent-count systems. Our work complements and extends these efforts by addressing fleet-scale modeling across thousands of agents in industrial settings.
@@ 4 Findings
  • Dynamic time warping (DTW) distance [42] across multiple dimensions: robot position and speed, state, and timing of load and unload events. Its units are the same as the underlying measurement, e.g., the DTW error for position is in meters, and measures the average distance between the true and estimated trajectory after the optimal temporal alignment is found.
@@
Based on these results, we conducted scaling experiments for the RC and GF architectures to extrapolate how more data and larger models using these architectures will improve performance. These experiments are modeled after scaling experiments in language modeling [31, 32], which have been used [34] to scale to extremely large models and datasets using architectural decisions made from smaller models.
@@ 5.1 Model Framework
This is similar to the decision transformer approach [43], but since we are in a behavior cloning setting our sequence alternates between state and action embeddings without returns-to-go. Note that since the action vocabulary is discrete, and the model outputs a probability distribution over the actions.
@@ 6 Robot-floor Model
Isolating the $N$-token robot stream $\mathbf{R}_{t}$ from the much longer $M$-token context stream $\mathbf{X}_{t}$ helps reduce quadratic self-attention cost from $O(M^{2})$ to $O(N^{2})$ (two orders of magnitude on a typical floor). Robots still access the full floor snapshot through a single cross-attention pass of cost $O(NM)$, which also injects rich relational features such as edge distance or path cost [44, 45]. This enables global context where all robots can, in principle, attend to every object on the warehouse floor (inventory, other robots, vertices and their connectivity), enabling long-range contextual understanding of fleet behavior. The context encoder can append new modalities (e.g., no-go zones) to $\mathbf{X}_{t}$ without touching the robot decoder, and the same model generalizes to arbitrary robot counts or floor sizes simply by changing $N$ and $M$.
@@ 6.1 Model Framework
The RF model treats next-action prediction as a sequence transduction task, mapping an input sequence of robot states to an output sequence of actions: $(\mathbf{r}_{1},\ldots,\mathbf{r}_{N})\mapsto(\mathbf{a}_{1},\ldots,\mathbf{a}_{N})$. To learn this mapping, interactions among robot and context tokens are modeled by two attention mechanisms. Self-attention among all robot tokens allows the model to learn robot-contextualized representations $p(\mathbf{r}_{i}^{\prime}|\mathbf{r}_{1},\ldots,\mathbf{r}_{N})$ that are predictive of next action (i.e., “how robots influence each other’s motion”). Cross-attention allows robot tokens to attend to the floor context around them to produce floor-contextualized representations $p(\mathbf{r}_{i}^{\prime}|\mathbf{x}_{1},\ldots,\mathbf{x}_{M})$ that are predictive of next action (i.e., “how the structure of the floor environment conditions next action probabilities”). The use of cross-attention allows very large contextual attention independent of the decoded sequence length [46].
@@ 7 Image-Floor Model
Within the DeepFleet design matrix it occupies the quadrant of fixed-time, synchronous predictors that enjoy full spatial context while modeling time autoregressively. Like video prediction models [47, 48], the IF model consumes a window of past floor frames and extrapolates the next frame; repeating this step yields a multi-step forecast. This dense, image-level strategy complements the sparse view of the RC model and the tokenized views of the RF and GF models by enabling direct pixel-wise reasoning over topology, traffic density, and local congestion patterns that emerge only when the floor is viewed as a whole. Because every robot is updated simultaneously, the model can exploit GPU-friendly batched convolutions and achieve constant latency—independent of fleet size—during both training and inference.

### v2

@@ 1 Introduction
The discovery that scaling up deep neural networks and their training data can unlock unprecedented capabilities in natural language understanding [brown2020language], speech recognition [zhang2022bigssl], vision [radford2021learning], and reasoning [openai2024o1] has recently led to efforts to apply the same data-driven scaling approach to robotic systems [WaymoScaling2025, black2024pi_0, oneill2024open, ghosh2024octo]. Inspired by these efforts, we explore the design space of multi-agent foundation models to support the development of the next generation of robot fleet coordination algorithms. Just as foundation models pretrained on simple tasks such as masked text prediction can be adapted to applications via supervised fine-tuning, reinforcement learning, knowledge distillation, or in-context learning, our multi-agent fleet models are designed to serve as a foundation for many applications in our warehouses, such as congestion forecasting, adaptive routing, and proactive rescheduling.
@@ 2 Related Work
Significant work has been dedicated to modeling multi-agent interactions in trajectory forecasting, particularly in domains such as pedestrian motion prediction, autonomous driving, and social navigation. Early works such as Social-LSTM [alahi2016social] and Social-GAN [gupta2018social] captured inter-agent dynamics via recurrent neural networks with interaction-aware pooling. Transformer-based architectures [giuliari2021transformer, salzmann2020trajectronpp], variational models [mangalam2021goals, tang2021collaborative], and hybrid methods combining GNNs with probabilistic forecasting [salzmann2020trajectron++, wang2024simple] have since improved generalization and multimodal prediction capabilities. However, these models are primarily applied to small-scale environments with limited agent counts and are not designed to scale to the thousands of agents encountered in warehouse scenarios.
@@
Research in warehouse robotics has historically focused on task assignment, multi-agent pathfinding (MAPF) [stern2019multi], and traffic management. Until recently, the most successful techniques have been heuristic-based methods such as conflict-based search (CBS) [boyarski2015icbs, li_eecbs_2021], prioritized planning [silver_cooperative_2021, ma_searching_2019], MAPF-LNS [li_anytime_2021], priority inheritance with backtracking (PIBT) [okumura_priority_2022], LaCAM3 [okumura_engineering_2024] and their variants — these often achieve strong performance with abstract robot models but struggle to scale on realistic robot models that account for dynamics and physical constraints [yan_multi-agent_2025, zhang_efficient_2023].
Learning-based techniques have also emerged to address path planning and task reallocation [fung2019coordinating, choudhury2022dynamic, sartoretti_primal_2019, liu_mapper_2020, damani_primal_2021, wang_scrimp_2023, wang_mobile_2020, tang_railgun_2025] (see [alkazzi_comprehensive_2024] for a review of learning-based MAPF through 2024), often combining imitation learning (IL) to learn behavior from heuristic MAPF solvers with reinforcement learning (RL) to learn improved behavior.
In 2024, Yan & Wu [yan_neural_2023] and Veerapaneni et al. [veerapaneni_improving_2024] independently proposed using learned action prediction models to guide heuristic search techniques; meanwhile, Zhang et al. proposed providing guidance with a model that generates graph edge weights [zhang_guidance_2024]. These learned guidance mechanisms steer the search by replacing myopic heuristics with data-driven ones, leading to faster runtimes and better solutions.
This combination of search and learning is emerging as a powerful tool in MAPF, propelling state of the art techniques such as SILLM [jiang_deploying_2025], SSIL [veerapaneni_work_2024], online GGO [zang_online_2025], and LaGAT [jain_graph_2025]. Notably, LaGAT [jain_graph_2025] demonstrates that even state-of-the-art learning-only methods like MAPF-GPT [andreychuk2025mapf] can be integrated into search frameworks, though search efficiency varies with model complexity.
The unique scale of both our warehouses and the data we collect from them allows us to truly test the limits of learning-only methods; in Section 4 we describe scaling laws over several orders of magnitude of model and training dataset size.
@@
Recent work in spatiotemporal modeling has produced powerful neural representations for structured environments. TrajectoryCNN [liu2020trajectorycnn], Scene-Transformer [ngiam2021scene], and ST-GNNs [bui2022spatial] illustrate the utility of combining convolutional or graph-based spatial reasoning with temporal sequence modeling. In robotics, structured state-space models (SSMs) [gu2023mamba, bhirangi2024hierarchical] and graph-based encodings [dik2024graph] have improved data efficiency and temporal coherence. Notably, GraphCast [lam2023learning] and GenCast [price2023gencast] have demonstrated the effectiveness of graph-based and generative models for large-scale spatiotemporal prediction, respectively excelling at deterministic and probabilistic forecasting in highly complex systems. Nonetheless, these methods often lack scalability, task generalization, or the integration of diverse spatial and temporal contexts required for warehouse-scale fleet modeling.
@@
The rise of foundation models was enabled by empirical studies [kaplan2020scaling] that demonstrated that language model performance followed predictable scaling laws with respect to compute, model size, and data. After further refinement [hoffmann2022training], these principles have been used to inform the scaling of large multi-modal models such as PaLM [chowdhery2022palm] and LLaMA [grattafiori_llama_2024].
@@
Foundation models are also achieving state-of-the-art capabilities in robotics. Models such as RT-1 [brohan2022rt], RT-2 [zitkovich2023rt], VIMA [jiang2022vima], and PerAct [shridhar2023perceiver] have demonstrated strong performance in manipulation and visuomotor policy learning. Flamingo [alayrac2022flamingo] has enabled visual reasoning for embodied agents, while $\pi_{0}$ [black2024pi_0], Octo [ghosh2024octo], and Helix [helix2025vision] have extended this foundation model approach toward multi-task, multi-embodiment settings. DeepMind’s work on robotics foundation models and scalable policy architectures [reed2022generalist] highlights the increasing effort to unify policy learning, perception, and reasoning in a single framework. These models emphasize generalization, multi-task capability, and real-world robustness—but are still primarily focused on manipulation or low-agent-count systems. Our work complements and extends these efforts by addressing fleet-scale modeling across thousands of agents in industrial settings.
@@ 4 Findings
  • Dynamic time warping (DTW) distance [muller2007dynamic] across multiple dimensions: robot position and speed, state, and timing of load and unload events. Its units are the same as the underlying measurement, e.g., the DTW error for position is in meters, and measures the average distance between the true and estimated trajectory after the optimal temporal alignment is found.
@@
Based on these results, we conducted scaling experiments for the RC and GF architectures to extrapolate how more data and larger models using these architectures will improve performance. These experiments are modeled after scaling experiments in language modeling [kaplan2020scaling, hoffmann2022training], which have been used [grattafiori_llama_2024] to scale to extremely large models and datasets using architectural decisions made from smaller models.
@@ 5.1 Model Framework
This is similar to the decision transformer approach [chen2021decision], but since we are in a behavior cloning setting our sequence alternates between state and action embeddings without returns-to-go. Note that since the action vocabulary is discrete, and the model outputs a probability distribution over the actions.
@@ 6 Robot-floor Model
Isolating the $N$-token robot stream $\mathbf{R}_{t}$ from the much longer $M$-token context stream $\mathbf{X}_{t}$ helps reduce quadratic self-attention cost from $O(M^{2})$ to $O(N^{2})$ (two orders of magnitude on a typical floor). Robots still access the full floor snapshot through a single cross-attention pass of cost $O(NM)$, which also injects rich relational features such as edge distance or path cost [shaw2018, press2021alibi]. This enables global context where all robots can, in principle, attend to every object on the warehouse floor (inventory, other robots, vertices and their connectivity), enabling long-range contextual understanding of fleet behavior. The context encoder can append new modalities (e.g., no-go zones) to $\mathbf{X}_{t}$ without touching the robot decoder, and the same model generalizes to arbitrary robot counts or floor sizes simply by changing $N$ and $M$.
@@ 6.1 Model Framework
The RF model treats next-action prediction as a sequence transduction task, mapping an input sequence of robot states to an output sequence of actions: $(\mathbf{r}_{1},\ldots,\mathbf{r}_{N})\mapsto(\mathbf{a}_{1},\ldots,\mathbf{a}_{N})$. To learn this mapping, interactions among robot and context tokens are modeled by two attention mechanisms. Self-attention among all robot tokens allows the model to learn robot-contextualized representations $p(\mathbf{r}_{i}^{\prime}|\mathbf{r}_{1},\ldots,\mathbf{r}_{N})$ that are predictive of next action (i.e., “how robots influence each other’s motion”). Cross-attention allows robot tokens to attend to the floor context around them to produce floor-contextualized representations $p(\mathbf{r}_{i}^{\prime}|\mathbf{x}_{1},\ldots,\mathbf{x}_{M})$ that are predictive of next action (i.e., “how the structure of the floor environment conditions next action probabilities”). The use of cross-attention allows very large contextual attention independent of the decoded sequence length [jaegle2021perceivergeneralperceptioniterative].
@@ 7 Image-Floor Model
Within the DeepFleet design matrix it occupies the quadrant of fixed-time, synchronous predictors that enjoy full spatial context while modeling time autoregressively. Like video prediction models [xing2024survey, zhou2024survey], the IF model consumes a window of past floor frames and extrapolates the next frame; repeating this step yields a multi-step forecast. This dense, image-level strategy complements the sparse view of the RC model and the tokenized views of the RF and GF models by enabling direct pixel-wise reasoning over topology, traffic density, and local congestion patterns that emerge only when the floor is viewed as a whole. Because every robot is updated simultaneously, the model can exploit GPU-friendly batched convolutions and achieve constant latency—independent of fleet size—during both training and inference.
@@ 9 Conclusions
10 List of Contributors
Please address correspondence to deepfleet@amazon.com.
Scaling experiments and writing lead
Ameya Agaskar
Sriram Siva
DeepFleet model design and development
William Pickering (Robot-Centric model)
Kyle O’Brien (Robot-Floor Cross Attention model)
Charles Kekeh (Image-Based Floor-Centric model)
Sriram Siva (Graph-Based Floor-Centric model)
Comparative empirical evaluation
Ang Li
Brianna Gallo Sarker
Alicia Chua
Data pipelines and infrastructure
Mayur Nemade
Charun Thattai
Jiaming Di
Isaac Iyengar
Ramya Dharoor
Dino Kirouani
Project management
Jimmy Erskine
Tamir Hegazy
Scientific guidance
Scott Niekum (Amazon Scholar)
Usman A. Khan (Amazon Scholar)
PI and research program lead
Federico Pecora
Joseph W. Durham
Acknowledgments
The authors wish to thank Zhe Chen for helpful discussions on multi-agent pathfinding (MAPF).